You exert an upward force of 6.5N on 1 pound of sugar. Consider the FG on sugar. What is the net force on the sugar?


1
Expert's answer
2022-03-07T10:19:10-0500

The net force

"F_{net}=F-mg\\
=\rm 6.5\: N-0.45\: kg*9.8\: N/kg=2.1\: N"